# Language Model Fine-Tuning on Scaled Survey Data for Predicting Distributions of Public Opinions

### v1

@@ 1 Introduction
Training on public opinion survey data has remained under-explored due to the limited availability of structured survey datasets. To this end, we curate and release SubPOP (Subpopulation-level Public Opinion Prediction), a dataset of 70K subpopulation-response distribution pairs (6.5×6.5$\times$6.5 × larger compared to previous datasets). We show that fine-tuning LLMs on SubPOP significantly improves the distributional match between LLM generated and human responses, and improvements are consistent across subpopulations of varying sizes. Additionally, the improvement generalizes to unseen subpopulations, survey waves (topics), and survey families, i.e. surveys administered by different institutions. Such broad generalization is particularly critical for real-world public opinions research, where practitioners are most in need of synthetic data for survey questions or subpopulations (or both) that they have not tested before.
@@
  • We show that training LLMs on response distributions from survey data significantly improves their ability to predict the opinions of subpopulations, reducing the Wasserstein distance between LLM and human distributions by 32-46% compared to top-performing baselines. (Section 4.2)
  • We show that the performance of the fine-tuned LLMs strongly generalizes to out-of-distribution data, including unseen subpopulations, new survey waves, and different survey families. (Section 4.2 and Section 4.3)
  • We release SubPOP, a curated and pre-processed dataset of public opinion survey results that is 6.5×6.5$\times$6.5 × larger than existing datasets, enabling fine-tuning at scale.
@@ 3.1 Fine-tuning LLMs on Human Response Distributions
Our goal is to fine-tune an LLM to predict the distribution of responses for a multiple-choice question, conditioned on descriptions of a human subpopulation we want to simulate, typically a specific demographic, socioeconomic, or ideological subgroup. Consider the example in Figure 2: the question asks, “What do you think the chances are these days that a woman won’t get a job or promotion while an equally or less qualified man gets one instead?” The available responses are: A. Very likely, B. Somewhat likely, C. Not very likely, D. Very unlikely, and E. Refused. In this case, the LLM will output a probability for each of the tokens corresponding to the choices A through E, thereby generating a complete response distribution that we aim to align with the true distribution observed in survey data.
@@
where $D_{\mathrm{KL}}$ denotes the KL divergence. In the example shown in Figure 2, the model is trained to reduce the KL divergence between the target (survey-based) distribution over {$A,B,C,D,E$} and its predicted distribution for the subpopulation living in the Southern United States.
@@ 3.3 Evaluation Metric
We use Wasserstein distance (WD) to quantify how closely the model’s predicted opinion distribution matches human survey data Santurkar et al. (2023); Moon et al. (2024); Meister et al. (2024); Zhao et al. (2023). Formally, for a group $g$ representing some subpopulation and a question $q$ WD is defined as $\mathcal{WD}_{\theta}(q,g)=\mathcal{WD}(p_{H}(\mathcal{A}_{q}|q,g),p_{\theta}(\mathcal{A}_{q}|q,g))$ (see formula in Appendix B). Since WD is computed over ordinal values, we map the categorical answer options to numbers, such as mapping “Very likely” to 1, “Likely” to 2, and so on.
Some prior work utilizes one-hot accuracy Feng et al. (2024); Li et al. (2023) as an evaluation metric. However, one-hot accuracy only verifies whether the top-predicted choice matches the top human response, thereby discarding distributional information. In contrast, WD accounts for partial overlaps among the categories and reflects the ‘cost’ of shifting probability mass, providing a more nuanced assessment of distribution discrepancy. Consider the example question provided in Figure 2, where the human response distribution indicates that option B (“Somewhat likely”) is the most probable. Now consider two cases in which the model incorrectly predicts the top choice. In the first case, the model assigns a high probability to option A (“Very likely”), while in the second case, it assigns a high probability to option D (“Very unlikely”). Although one-hot accuracy would treat both predictions equally as errors, WD differentiates between them by accounting for the ordinal relationship among the options, penalizing the second prediction more heavily for its larger deviation from the true distribution.
@@ 4 Experiments
[…]
@@ 4.1 Bounds of WD and Baselines
We compare our approach with two baseline methods: prompting and Modular Pluralism Feng et al. (2024). For prompting, we consider both zero-shot and few-shot methods. In zero-shot prompting, we steer the LLM using demographic prompt formats. Specifically, we employ three different formats following Santurkar et al. (2023): QA, BIO, and PORTRAY. For instance, to condition the LLM to a person living in the South of the US, the QA format uses a question-answer format as illustrated in Figure 2; the BIO format conditions the model with a first-person narrative such as “I currently reside in the South."; and the PORTRAY format uses a third-person narrative like “Answer the following question as if you currently reside in the South.".
@@
Modular pluralism Feng et al. (2024) fine-tunes multiple LLMs on distinct datasets to capture the viewpoints of different communities Feng et al. (2023). For a given question, each fine-tuned LLM generates an opinion that reflects the perspective of the community it represents, and a separate black-box LLM aggregates these outputs to produce the final distributional response. Detailed implementation of the lower/upper bounds and the baselines is provided in Appendix D.
@@ 4.2 Generalization to Unseen Topics and Survey Families
In this section, we assess the ability of our fine-tuned LLMs to generalize to unseen data—both in terms of new topics and entirely different survey families. To evaluate these aspects, we use OpinionQA to measure generalization to unseen topics, and SubPOP-Eval to test generalization to a different survey family. We fine-tune four LLMs (Llama-2-7B, Llama-2-13B, Mistral-7B, and Llama-3-70B) on SubPOP-Train. We opt for pretrained LLMs rather than instruction-following models, as previous work has shown that pretrained models perform better on this task Moon et al. (2024). A detailed comparison between these model types is provided in Section C.2.
@@
Table 1 reports the average WD metrics computed over all demographic groups and survey questions, comparing our fine-tuned models against various baseline approaches. Our experiments show that fine-tuning on SubPOP-Train significantly outperforms all other methods, yielding a 32–46% reduction in WD on OpinionQA and a 39–42% reduction on SubPOP-Eval compared to the best baselines. Notably, SubPOP-Train is based on ATP data, while SubPOP-Eval is derived from GSS surveys—two distinct survey families that can differ in respondent pools, calibration techniques, and other methodological factors, leading to non-trivial distribution shifts despite both being representative of the US population. Furthermore, our fine-grained analyses at the wave level (see Appendix E) confirm that these trends persist even at more detailed levels of evaluation.
@@ 4.3 Generalization across Target Subpopulations
Figure 3: Per-group evaluation performance of our model Llama-2-7B-SubPOP-FT (red lines) on OpinionQA. For comparison, the results from zero-shot QA prompting (black lines) and the lower bound (blue lines) are presented. We observe that the relative improvement, measuring how much of the gap between zero-shot prompting and the lower bound has been closed, remains consistent across subpopulations. Shaded blue regions represent the 95% confidence interval of the lower-bound estimation for each group. Per-group results for other models (Table 9) and the results on SubPOP evaluation set (Table 10) are available in Appendix E.
@@
Figure 3 shows the per-group WD on the OpinionQA evaluation for Llama-2-7B, comparing our fine-tuning approach with zero-shot prompting and the empirical WD lower bound. To evaluate the consistency of performance gains, we calculate the relative improvement for each subpopulation as how much of the gap between zero-shot prompting and the empirical lower bound is reduced after fine-tuning. This measure allows us to account for varying lower bounds across subpopulations: since some groups have fewer respondents, there is greater uncertainty in their reported distribution in the survey data and greater variance between the original sample and bootstrap samples.
@@
As shown in Table 2, our model achieves a strong reduction in WD even for unseen subpopulations, indicating that the model can be steered by demographic prompts beyond the seen subpopulations in training. Interestingly, although SubPOP-Train does not contain any data with opinion distributions of particular age groups (e.g. subjects of age 18-29 or those of age 65+), the average relative improvement is 44.7%, which is compatible with the average relative improvement for seen subpopulations. We provide results for other unseen groups in Table 7 of Section C.3 (average relative improvement of 43.1% with a standard deviation of 6.7%).
@@
Formally, we define the intergroup disagreement between a target group $g_{t}$ and a source group $g_{s}$ as $\mathcal{WD}(p_{H}(\mathcal{A}_{q}\mid q,g_{t}),p_{H}(\mathcal{A}_{q}\mid q,g_{s}))$ averaged over evaluation questions. In human responses (left of Figure 4), the disagreement shows the pattern of locality: increases as the disparity in education levels between two groups grows. We extend this notion to compare the human distribution from the target group $g_{t}$ with the LLM-predicted distribution when the source group $g_{s}$ is specified in the prompt, $\mathcal{WD}(p_{H}(\mathcal{A}_{q}\mid q,g_{t}),p_{\theta}(\mathcal{A}_{q}\mid q,g_{s}))$. If the model truly incorporates subpopulation information from the prompt, its intergroup disagreement pattern should mirror that of the human data.
Zero-shot prompting with the base model (right of Figure 4) does not exhibit the locality pattern seen in the human data, indicating that it cannot be steered by subpopulation labels. In contrast, the fine-tuned model (middle of Figure 4) reproduces a pattern resembling the human-human case, even though it was trained on only two education groups (“less than high school” and “college graduate/some postgrad”) and the other four groups were unseen. This result demonstrates that our fine-tuned model not only learns to condition on subpopulation information but also generalizes to subpopulations unseen during fine-tuning. We provide the intergroup disagreement for other traits in Section C.3.
@@ 4.4 Effect of Scaling the Dataset
In this section, we examine performance scales with training dataset size. We randomly sample subsets containing 25%, 50%, 75%, and 87.5% of the full SubPOP training set and evaluate three models—Llama-2-7B, Llama-2-13B, and Mistral-7B—on OpinionQA. As shown in Figure 5, we observe diminishing marginal returns, as is typical with fine-tuning; for example, after training on a random 25%, the models reach 72%-78% of the total improvement they achieve after fine-tuning on all of SubPOP-train. However, the performance does not entirely plateau. Instead, it continues to improve as we further increase the training data from 25% to 100%. We fit linear trend lines (dotted in Figure 5) to the results and observe that the slopes are similar for each model. This suggests that the rate of improvement—reflected by the slope in the power-law relationship—is intrinsic to the data and task rather than to the specific model architecture. In other words, LLMs exhibit comparable data efficiency, with performance gains that are fundamentally tied to dataset size rather than model-specific factors.
Using these trend lines, we can estimate the amount of fine-tuning data required to reach a target performance. For instance, we estimate that fine-tuning Mistral-7B on a dataset 25 times larger than the current SubPOP training set would yield a WD value of 0.07, which is much closer to the empirical lower bound of 0.031 reported in Table 1. This result underscores the critical importance of collecting more high-quality data, as increased dataset size can drive significant improvements in model performance.
@@ A.2 OpinionQA
Subpopulations we employ are listed in Table 3. This set of groups is adopted for several small-scale analysis Santurkar et al. (2023); Zhao et al. (2023); Kim and Yang (2024). We note that our approach is not limited to a specific number of groups and data is available for small or fine-grained demographic subpopulations.
@@ A.3 SubPOP-Train
We gather additional data from the American Trends Panel, specifically collecting 53 waves from Wave 61 to 132. There are 62 waves from Wave 61 - 132, however, some waves have missing demographic or ideology information (for example, wave 63 does not contain political ideology information) or the data is not available hence removed during the curation process. To refine the dataset, we exclude questions that meet the following criteria: those with more than 10 response options, redacted response data, or dependencies on prior questions (e.g., assessing political strength). For the remaining questions, we use GPT-4o to refine their wording, ensuring they are well-suited for prompting the language models while making minimal modifications. In Figure 6 we provide a few-shot prompt for question refinement.
@@
In Figure 7, we visualize the embeddings of the question texts (projected to 2-dimensions using t-SNE) from OpinionQA compared to SubPOP-Train and SubPOP-Eval. The visualization shows how much larger our dataset is than OpinionQA (6.5$\times$), along with the expanded coverage of our dataset into semantic areas untouched by OpinionQA. The embeddings also reveal the distribution shift from ATP questions to GSS questions: while the ATP and GSS questions overlap in embedding space, the GSS question appear as small clusters, not evenly distributed over the ATP questions. In Table 4, we list each ATP wave in SubPOP-Train and OpinionQA, along with its number of questions and wave topic(s), as defined by ATP. [^1] ATP wave topics and time periods are defined at https://www.pewresearch.org/american-trends-panel-datasets/. The table indicates which topics are new in SubPOP-Train compared to OpinionQA, indicating the expanded coverage of our dataset, along with which topics remain unseen in OpinionQA, which we can use to test LLMs fine-tuned on SubPOP-Train for generalization.
@@ A.4 SubPOP-Eval
To further evaluate the out-of-distribution generalization ability of our fine-tuned models, we subsample 133 questions from the GSS 2022 dataset Davern et al. (2024). We apply the same selection criteria as outlined in  Section A.3, excluding questions that are redacted, conditioned on prior questions, inferable directly from the group information, derived from a set of questions, or those with more than 10 response options.

### v2

@@ 1 Introduction
Training on public opinion survey data has remained under-explored due to the limited availability of structured survey datasets. To this end, we curate and release SubPOP (Subpopulation-level Public Opinion Prediction), a dataset of 70K subpopulation-response distribution pairs ($6.5\times$ larger compared to previous datasets). We show that fine-tuning LLMs on SubPOP significantly improves the distributional match between LLM generated and human responses, and improvements are consistent across subpopulations of varying sizes. Additionally, the improvement generalizes to unseen subpopulations, survey waves (topics), and survey families, i.e. surveys administered by different institutions. Such broad generalization is particularly critical for real-world public opinions research, where practitioners are most in need of synthetic data for survey questions or subpopulations (or both) that they have not tested before.
@@
  • We show that training LLMs on response distributions from survey data significantly improves their ability to predict the opinions of subpopulations, reducing the Wasserstein distance between LLM and human distributions by 32-46% compared to top-performing baselines. (Section˜4.2)
  • We show that the performance of the fine-tuned LLMs strongly generalizes to out-of-distribution data, including unseen subpopulations, new survey waves, and different survey families. (Section˜4.2 and Section˜4.3)
  • We release SubPOP, a curated and pre-processed dataset of public opinion survey results that is $6.5\times$ larger than existing datasets, enabling fine-tuning at scale.
@@ 3.1 Fine-tuning LLMs on Human Response Distributions
Our goal is to fine-tune an LLM to predict the distribution of responses for a multiple-choice question, conditioned on descriptions of a human subpopulation we want to simulate, typically a specific demographic, socioeconomic, or ideological subgroup. Consider the example in Figure˜2: the question asks, “What do you think the chances are these days that a woman won’t get a job or promotion while an equally or less qualified man gets one instead?” The available responses are: A. Very likely, B. Somewhat likely, C. Not very likely, D. Very unlikely, and E. Refused. In this case, the LLM will output a probability for each of the tokens corresponding to the choices A through E, thereby generating a complete response distribution that we aim to align with the true distribution observed in survey data.
@@
where $D_{\mathrm{KL}}$ denotes the KL divergence. In the example shown in Figure˜2, the model is trained to reduce the KL divergence between the target (survey-based) distribution over {A,B,C,D,EA,B,C,D,E} and its predicted distribution for the subpopulation living in the Southern United States.
@@ 3.3 Evaluation Metric
We use Wasserstein distance (WD) to quantify how closely the model’s predicted opinion distribution matches human survey data Santurkar et al. (2023); Moon et al. (2024); Meister et al. (2024); Zhao et al. (2023). Formally, for a group $g$ representing some subpopulation and a question $q$ WD is defined as $\mathcal{WD}_{\theta}(q,g)=\mathcal{WD}(p_{H}(\mathcal{A}_{q}|q,g),p_{\theta}(\mathcal{A}_{q}|q,g))$ (see formula in Appendix˜B). Since WD is computed over ordinal values, we map the categorical answer options to numbers, such as mapping “Very likely” to 1, “Likely” to 2, and so on.
Some prior work utilizes one-hot accuracy Feng et al. (2024); Li et al. (2023) as an evaluation metric. However, one-hot accuracy only verifies whether the top-predicted choice matches the top human response, thereby discarding distributional information. In contrast, WD accounts for partial overlaps among the categories and reflects the ‘cost’ of shifting probability mass, providing a more nuanced assessment of distribution discrepancy. Consider the example question provided in Figure˜2, where the human response distribution indicates that option B (“Somewhat likely”) is the most probable. Now consider two cases in which the model incorrectly predicts the top choice. In the first case, the model assigns a high probability to option A (“Very likely”), while in the second case, it assigns a high probability to option D (“Very unlikely”). Although one-hot accuracy would treat both predictions equally as errors, WD differentiates between them by accounting for the ordinal relationship among the options, penalizing the second prediction more heavily for its larger deviation from the true distribution.
@@ 4 Experiments
Table 1: Evaluation on OpinionQA and the SubPOP evaluation set (SubPOP-Eval) for 22 subpopulations following  Santurkar et al. (2023). We compute the WD by averaging over all questions and subpopulations. Lower and upper bounds of performance give guidance on how each method performs. For Modular Pluralism, we provide an error rate of one-hot prediction ($\dagger$}) (Section˜3.3) which was used in the original paper.
@@ 4.1 Bounds of WD and Baselines
We compare our approach with two baseline methods: prompting and Modular Pluralism Feng et al. (2024). For prompting, we consider both zero-shot and few-shot methods. In zero-shot prompting, we steer the LLM using demographic prompt formats. Specifically, we employ three different formats following Santurkar et al. (2023): QA, BIO, and PORTRAY. For instance, to condition the LLM to a person living in the South of the US, the QA format uses a question-answer format as illustrated in Figure˜2; the BIO format conditions the model with a first-person narrative such as “I currently reside in the South."; and the PORTRAY format uses a third-person narrative like “Answer the following question as if you currently reside in the South.".
@@
Modular pluralism Feng et al. (2024) fine-tunes multiple LLMs on distinct datasets to capture the viewpoints of different communities Feng et al. (2023). For a given question, each fine-tuned LLM generates an opinion that reflects the perspective of the community it represents, and a separate black-box LLM aggregates these outputs to produce the final distributional response. Detailed implementation of the lower/upper bounds and the baselines is provided in Appendix˜D.
@@ 4.2 Generalization to Unseen Topics and Survey Families
In this section, we assess the ability of our fine-tuned LLMs to generalize to unseen data—both in terms of new topics and entirely different survey families. To evaluate these aspects, we use OpinionQA to measure generalization to unseen topics, and SubPOP-Eval to test generalization to a different survey family. We fine-tune four LLMs (Llama-2-7B, Llama-2-13B, Mistral-7B, and Llama-3-70B) on SubPOP-Train. We opt for pretrained LLMs rather than instruction-following models, as previous work has shown that pretrained models perform better on this task Moon et al. (2024). A detailed comparison between these model types is provided in Section˜C.2.
@@
Table˜1 reports the average WD metrics computed over all demographic groups and survey questions, comparing our fine-tuned models against various baseline approaches. Our experiments show that fine-tuning on SubPOP-Train significantly outperforms all other methods, yielding a 32–46% reduction in WD on OpinionQA and a 39–42% reduction on SubPOP-Eval compared to the best baselines. Notably, SubPOP-Train is based on ATP data, while SubPOP-Eval is derived from GSS surveys—two distinct survey families that can differ in respondent pools, calibration techniques, and other methodological factors, leading to non-trivial distribution shifts despite both being representative of the US population. Furthermore, our fine-grained analyses at the wave level (see Appendix˜E) confirm that these trends persist even at more detailed levels of evaluation.
@@ 4.3 Generalization across Target Subpopulations
Figure 3: Per-group evaluation performance of our model Llama-2-7B-SubPOP-FT (red lines) on OpinionQA. For comparison, the results from zero-shot QA prompting (black lines) and the lower bound (blue lines) are presented. We observe that the relative improvement, measuring how much of the gap between zero-shot prompting and the lower bound has been closed, remains consistent across subpopulations. Shaded blue regions represent the 95% confidence interval of the lower-bound estimation for each group. Per-group results for other models (Table˜9) and the results on SubPOP evaluation set (Table˜10) are available in Appendix˜E.
@@
Figure˜3 shows the per-group WD on the OpinionQA evaluation for Llama-2-7B, comparing our fine-tuning approach with zero-shot prompting and the empirical WD lower bound. To evaluate the consistency of performance gains, we calculate the relative improvement for each subpopulation as how much of the gap between zero-shot prompting and the empirical lower bound is reduced after fine-tuning. This measure allows us to account for varying lower bounds across subpopulations: since some groups have fewer respondents, there is greater uncertainty in their reported distribution in the survey data and greater variance between the original sample and bootstrap samples.
@@
As shown in Table 2, our model achieves a strong reduction in WD even for unseen subpopulations, indicating that the model can be steered by demographic prompts beyond the seen subpopulations in training. Interestingly, although SubPOP-Train does not contain any data with opinion distributions of particular age groups (e.g. subjects of age 18-29 or those of age 65+), the average relative improvement is 44.7%, which is compatible with the average relative improvement for seen subpopulations. We provide results for other unseen groups in Table˜7 of Section˜C.3 (average relative improvement of 43.1% with a standard deviation of 6.7%).
@@
Formally, we define the intergroup disagreement between a target group $g_{t}$ and a source group $g_{s}$ as $\mathcal{WD}(p_{H}(\mathcal{A}_{q}\mid q,g_{t}),p_{H}(\mathcal{A}_{q}\mid q,g_{s}))$ averaged over evaluation questions. In human responses (left of Figure˜4), the disagreement shows the pattern of locality: increases as the disparity in education levels between two groups grows. We extend this notion to compare the human distribution from the target group $g_{t}$ with the LLM-predicted distribution when the source group $g_{s}$ is specified in the prompt, $\mathcal{WD}(p_{H}(\mathcal{A}_{q}\mid q,g_{t}),p_{\theta}(\mathcal{A}_{q}\mid q,g_{s}))$. If the model truly incorporates subpopulation information from the prompt, its intergroup disagreement pattern should mirror that of the human data.
Zero-shot prompting with the base model (right of Figure˜4) does not exhibit the locality pattern seen in the human data, indicating that it cannot be steered by subpopulation labels. In contrast, the fine-tuned model (middle of Figure˜4) reproduces a pattern resembling the human-human case, even though it was trained on only two education groups (“less than high school” and “college graduate/some postgrad”) and the other four groups were unseen. This result demonstrates that our fine-tuned model not only learns to condition on subpopulation information but also generalizes to subpopulations unseen during fine-tuning. We provide the intergroup disagreement for other traits in Section˜C.3.
@@ 4.4 Effect of Scaling the Dataset
In this section, we examine performance scales with training dataset size. We randomly sample subsets containing 25%, 50%, 75%, and 87.5% of the full SubPOP training set and evaluate three models—Llama-2-7B, Llama-2-13B, and Mistral-7B—on OpinionQA. As shown in Figure˜5, we observe diminishing marginal returns, as is typical with fine-tuning; for example, after training on a random 25%, the models reach 72%-78% of the total improvement they achieve after fine-tuning on all of SubPOP-train. However, the performance does not entirely plateau. Instead, it continues to improve as we further increase the training data from 25% to 100%. We fit linear trend lines (dotted in Figure˜5) to the results and observe that the slopes are similar for each model. This suggests that the rate of improvement—reflected by the slope in the power-law relationship—is intrinsic to the data and task rather than to the specific model architecture. In other words, LLMs exhibit comparable data efficiency, with performance gains that are fundamentally tied to dataset size rather than model-specific factors.
Using these trend lines, we can estimate the amount of fine-tuning data required to reach a target performance. For instance, we estimate that fine-tuning Mistral-7B on a dataset 25 times larger than the current SubPOP training set would yield a WD value of 0.07, which is much closer to the empirical lower bound of 0.031 reported in Table˜1. This result underscores the critical importance of collecting more high-quality data, as increased dataset size can drive significant improvements in model performance.
@@ A.2 OpinionQA
Subpopulations we employ are listed in Table˜3. This set of groups is adopted for several small-scale analysis Santurkar et al. (2023); Zhao et al. (2023); Kim and Yang (2024). We note that our approach is not limited to a specific number of groups and data is available for small or fine-grained demographic subpopulations.
@@ A.3 SubPOP-Train
We gather additional data from the American Trends Panel, specifically collecting 53 waves from Wave 61 to 132. There are 62 waves from Wave 61 - 132, however, some waves have missing demographic or ideology information (for example, wave 63 does not contain political ideology information) or the data is not available hence removed during the curation process. To refine the dataset, we exclude questions that meet the following criteria: those with more than 10 response options, redacted response data, or dependencies on prior questions (e.g., assessing political strength). For the remaining questions, we use GPT-4o to refine their wording, ensuring they are well-suited for prompting the language models while making minimal modifications. In Figure˜6 we provide a few-shot prompt for question refinement.
@@
[…]
@@ A.4 SubPOP-Eval
To further evaluate the out-of-distribution generalization ability of our fine-tuned models, we subsample 133 questions from the GSS 2022 dataset Davern et al. (2024). We apply the same selection criteria as outlined in  Section˜A.3, excluding questions that are redacted, conditioned on prior questions, inferable directly from the group information, derived from a set of questions, or those with more than 10 response options.